If; If f(x)={(3,5),(2,4),(1,7)} and h(x)={(3,2),(4,3),(1,6)}
(f+h)(1)=

Answer by jim_thompson5910(35256)

(f+h)(1)= f(1) + h(1) = 7 + 6 = 13


So (f+h)(1) = 13